If IL is 20 dB then sound find corresponding wave ? intensity of the
Answer #1

Intensity of the sound wave in dB = 10 log (I / Io)

log ( I / I o) = 20/10 = 2

I o = 10 -12 W / m2

I = e 2 x 10 -12  = 7.38 x 10 -12 W / m2